The site is rectangular with a permiter of 84m. Its length is 8m more than 4 times its width.
Work out the length and width of the rectangle you must present your answere with the corrct decimal place.

Answer by fractalier(6550)

Perimeter P = 2L + 2W = 84. We also have
L = 4W + 8
Now substitute that into the first equation and get
2(4W + 8) + 2W = 84
8W + 16 + 2W = 84
10W + 16 = 84
10W = 68
W = 6.8 m
L = 4(6.8) + 8 = 35.2 m